Prove that if A is a square matrix then AA^T and A + A are symmetric


1
Expert's answer
2021-05-23T20:07:02-0400

Let A be a square matrix.

Consider

"(AA^T)^T = (A^T)^TA^T= AA^T"

Hence "AA^T" is symmetric

Also, consider

"(A+A^T)^T = A^T + (A^T)^T = A^T + A = A+A^T"

Hence "A+A^T" is symmetric.